OpenStudy (kewlgeek555):

Which exponential function goes through the points (1, 8) and (4, 64)?

OpenStudy (ranga):

Two points are given for the function: (1, 8) and (4, 64) Put x = 1 and see which one gives you 8. And put x = 4 and see which one gives you 64. Both points should be satisfied and when you try both points will know which one does.

OpenStudy (kewlgeek555):

\[a. f(x)=4(2)^{x}\]\[b. f(x)=2(4)^{x}\]\[c. f(x)=4(2)^{-x}\]\[d. f(x)=2(4)^{-x}\]

OpenStudy (kewlgeek555):

Okay...will do.

OpenStudy (kewlgeek555):

1 & 16?

OpenStudy (kewlgeek555):

That sounds wrong...

OpenStudy (ranga):

Let us take a) f(x) = 4(2)^x put x = 1 y = 4(2)^1 = 8 So it satisfies (1,8) put x = 4 y = 4(2)^4 = 4 * 16 = 64. So it satisfies (4,64) So both points are on the curve and so the first function goes through both points. Try the other function choices too.

OpenStudy (kewlgeek555):

Ah. So it must be A, right? Since it satisfies both points... I'll do the other ones for you.

OpenStudy (ranga):

Yeah it is a). But it would be good to make sure the other functions don't satisfy both points.

OpenStudy (kewlgeek555):

\[f(x)=2(4)^{x}\]\[f(1)=2 \times 4\]\[f(1)=2 \times 4\]\[f(1)=8\]

OpenStudy (kewlgeek555):

\[f(x)=2(4)^{x}\]\[f(4)=2 \times 4^{4}\]\[f(4)=2 \times 256\]\[f(4) = 512\]

OpenStudy (kewlgeek555):

So B satisfies (1, 8), but not (4, 64).

OpenStudy (ranga):

yes. so it is not b.

OpenStudy (kewlgeek555):

Nerp.

OpenStudy (ranga):

c) and d) don't even satisfy the first point and so they are out too.

OpenStudy (kewlgeek555):

So it's A!

OpenStudy (kewlgeek555):

Can you help me on another one? I just need you to check if it is correct. thnx.